# Patterns of X-ray and -ray Flares: from Blazar to Maganetar and Sun

### v1

@@ 4 Discussion
The variability of the solar X-ray flares, the blazar -ray flares and the FRB-associated HXMT-HE XRB in SGR 1935+2154 all follow SHO model, yet they exhibit differences in values, PSDs, and characteristic timescales as shown in FigureĀ 5 and TableĀ 2.
@@
We aim to investigate the underlying mechanisms of blazar extreme -ray flares and XRBs in magnetar through a comparative framework, leveraging insights from solar X-ray flares fitting results and their associated physical origins. The driving mechanism of solar X-ray flares is widely attributed to magnetic reconnection, which converts stored magnetic energy into plasma heating and particle acceleration (Krucker etĀ al., 2008; Aschwanden, 2019). In the solar X-ray flare system and in the framework of magnetic reconnection, the timescale (4-7 s) likely corresponds to the energy injection phase, and the flat decay timescale (16-53 s) may reflect subsequent dissipative processes, such as plasma cooling and particle thermalization (e.g., Bian etĀ al., 2012; Aschwanden etĀ al., 2015). The specific dynamical behavior is intrinsically governed by the -factor of the system. Solar X-ray flare 3 exhibits different values compared to flares 1 and 2, potentially due to differences in magnetic topology or plasma environment (Benz, 2017).
@@
The discovery of the FRB-associated HXMT-HE XRB in SGR 1935+2154 suggests that a physical process in magnetar activity (such as magnetic reconnection or oscillation modes) simultaneously leads to the generation of both XRBs and FRBs (CHIME/FRB Collaboration etĀ al., 2020; Li etĀ al., 2021). The critically damped state of a system could be the condition triggering this combined event. In this case, the critically damped regime of this actual system corresponds to , which differs from the theoretical value of . In the comparative framework for the different types of extreme flares, the FRB-associated HXMT-HE XRB have similar values to that of solar flares 1 2, indicating similar dynamic behavior. The difference in characteristic timescale may be attributed to the difference in magnetic field strength. In addition, the ultra-short timescales of FRB-associated HXMT-HE XRB in SGR 1935+2154 (0.005 s) are consistent with Alfvnic crossing times in magnetar magnetospheres (Li & Beloborodov, 2015; Yuan etĀ al., 2020), supporting fast magnetic reconnection triggered by sudden magnetic field reconfiguration.
@@
For blazar -ray flares, the central values of exhibit an upward shift from 0.3, similar to that of solar X-ray flare 3, indicating that they may share certain common dynamic characteristics within their respective systems. A 13-hour quasi-periodic oscillation (QPO) signal detected across -ray flux, optical flux, and linear polarization during the flaring state of BL Lacertae has been reported by Jorstad etĀ al. (2022), in which case the values could lie in the range greater than 1. This result strongly supports an outburst mechanism, i.e., current-driven kink instabilities (e.g., Dong etĀ al., 2020), that modulates the reconnection process.
@@
The diversity in dynamical behaviors among solar X-ray flares, magnetar XRBs, and blazar flares likely stems from differences in magnetic field geometries (e.g., solar loops vs. relativistic jets) and particle acceleration efficiency. For instance, the 2ā5 day timescales of blazar -ray flares may correspond to the light-crossing time of compact emission regions, where it may be the magnetic reconnection in turbulent jets driving rapid particle acceleration. While the second-scale solar X-ray flares and millisecond-scale XRBs in magnetars may reflect energy release processes directly. These high-energy outburst events may share a fundamental reliance on magnetic energy dissipation, such as magnetic reconnection.
An ongoing issue is that, in the light curve fitting, although the ACF of the residual squares mostly falls within the white noise range, it does not randomly distribute around zero, particularly for the solar X-ray flare 3. Future research requires more accurate models and higher-quality data.

### v2

@@ 4 Discussion
The variability of the solar X-ray flares, the blazar -ray flares and the FRB-associated HXMT-HE XRB in SGR 1935+2154 all follow SHO model, yet they exhibit differences in values, PSDs, and characteristic timescales as shown in Figure 6 and Table 2.
@@
We aim to investigate the underlying mechanisms of flaring events from different astronomical objects through a comparative framework, leveraging insights from solar X-ray flares fitting results and their associated physical origins. The driving mechanism of solar X-ray flares is widely attributed to magnetic reconnection, which converts stored magnetic energy into plasma heating and particle acceleration (2008A&ARv..16..155K; 2019ASSL..458.....A). In the solar X-ray flare system and in the framework of magnetic reconnection, the timescale (4-7 s) likely corresponds to the energy injection phase, and the flat decay timescale (16-53 s) may reflect subsequent dissipative processes, such as plasma cooling and particle thermalization (e.g., 2012ApJ...754..103B; 2015ApJ...802...53A). These timescales should be interpreted with caution.
For solar flares, we note that the high-frequency broken timescale ( 4-7s) is broadly consistent with the transit time in compact coronal loops, suggesting a connection to rapid magnetohydrodynamics (MHD) energy transport or particle acceleration (2009SSRv..149..119N; 2018SSRv..214...45M). It also closely matches few-second QPP periods observed in hard X-ray emissions and predicted by MHD simulations (2020STP.....6a...3K; 2023BAAS...55c.181I). In contrast, the low-frequency broken timescale (16-53 s) overlaps with the most commonly reported QPP periods in GOES soft X-rays (10-40 s; 2020ApJ...895...50H) and some hard X-ray emissions (e.g., 2017MNRAS.471L...6L). This overlap between the timescales we obtained and previously reported QPP periods in solar flares suggests that a common underlying energy release mechanism, such as intermittent/repetitive magnetic reconnection, global MHD oscillatory modes (e.g., kink and sausage modes) in coronal structures, or an associated modulation process may be at work (2005psci.book.....A; 2009SSRv..149..119N; 2013A&A...558A..76R; 2016SoPh..291.3143V; 2021SoPh..296..188Z; 2023BAAS...55c.181I).
It is important to emphasize that although the PSD-derived characteristic timescales from our GP analysis overlap in magnitude with QPP oscillation timescales in solar flares, they are not identical. The latter represents explicit oscillatory signals (periodic timescale) that may arise from the underlying energy release dynamics, whereas the former reflect the underlying characteristic timescales of the energy release process itself, even in the absence of a dominant coherent oscillation. In other words, our analysis reveals the fundamental timescales of energy injection and dissipation that govern the system’s behavior, of which QPPs may be one particular manifestation. This provides a complementary perspective on the multi-scale nature of reconnection-driven energy release, but they do not uniquely determine the underlying mechanism. Future high-cadence imaging observations, multi-wavelength spectral diagnostics, and quantitative MHD modeling will be essential for distinguishing between these scenarios.
@@
The discovery of the FRB-associated HXMT-HE XRB in SGR 1935+2154 suggests that a physical process in magnetar activity (such as magnetic reconnection or oscillation modes) simultaneously leads to the generation of both XRBs and FRBs (2020Natur.587...54C; 2021NatAs...5..378L). The critically damped state of a system could be the condition triggering this combined event. In this case, the critically damped regime of this actual system corresponds to , which differs from the theoretical value of . In the comparative framework for the different types of extreme flares, the FRB-associated HXMT-HE XRB have similar values to that of solar flares 1 2, indicating similar dynamic behavior. The ultra-short timescales of FRB-associated HXMT-HE XRB in SGR 1935+2154 (0.005 s) are consistent with Alfvnic crossing times in magnetar magnetospheres (2015ApJ...815...25L; 2020ApJ...900L..21Y), supporting fast magnetic reconnection triggered by sudden magnetic field reconfiguration. This supports a similar energy release process between the two class outbursts. The difference in characteristic timescale may be attributed to the difference in magnetic field strength and emission scale.
@@
For blazar -ray flares, the central values of exhibit an upward shift from 0.3, similar to that of solar X-ray flare 3, indicating that they may share certain common dynamic characteristics within their respective systems. In addition, a 13-hour QPO signal detected across -ray flux, optical flux, and linear polarization during the flaring state of BL Lacertae (a blazar) has been reported by 2022Natur.609..265J, and it strongly support an outburst mechanism, i.e., current-driven kink instabilities (e.g., 2020MNRAS.494.1817D), that modulates the reconnection process.
@@
The diversity in dynamical behaviors among solar X-ray flares, magnetar XRBs, and blazar flares likely stems from differences in magnetic field geometries (e.g., solar loops vs. relativistic jets) and particle acceleration efficiency. For instance, the 2-5 day timescales of blazar -ray flares may correspond to the light-crossing time of compact emission regions, where it may be the magnetic reconnection in turbulent jets driving rapid particle acceleration. While the second-scale solar X-ray flares and millisecond-scale XRBs in magnetars may reflect energy release processes directly. These high-energy outburst events may share a fundamental reliance on magnetic energy dissipation, such as magnetic reconnection.
There are several ongoing issues. The first is that in the light curve fitting, although the ACF of the residual squares mostly falls within the white noise range, it does not randomly distribute around zero, particularly for the solar X-ray flare 3. Maybe it needs further analysis. A second issue is that the SHO model provides the statistically optimal fit under the criterion for the flare events analyzed in this study, but its derived physical interpretation must be treated with caution. The inferences about flare physics based on the SHO model’s Q-values and PSD structure should be regarded as exploratory at this stage. Confirmation of the SHO model’s physical superiority over simpler models and the universality of the inferred Q-values requires analysis of a larger and more diverse set of flare events, which we identify as a key direction for future work. Furthermore, Our analysis focuses on the near-peak phase (approximate impulsive phase) of the flare. When testing an expanded time interval to cover a larger portion of the light curve, we found that none of the SHO, DRW, Matrn models, or their combinations can adequately capture the variability across the extended time interval, indicating non-stationary data or the presence of multiple physical regimes. Alternative approaches such as the changepoints method may be required, which expresses a change from one kernel to another. We are currently developing this approach.